• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name friedman

    Log entry time 08:51:13 on March 11, 2012

    Entry number 365648

    keyword=owl shift summary

    00:22 - Spot++ looks fine.
    00:41 - Third arm coda restated.
    01:03 - We stop the beam to repolarize target. Tried negative polarization.
    01:34 - Beam back on.
    01:50 - IHWP is now IN.
    02:00 - Time changed to 03:00!!!
    03:02 - Runs 3071, 22156, 40303 wasn't stopped before IHWP changed.
    03:10 - Online DAQ ceashes.
    03:35 - Bleedtrough ceack required from MCC - no more than 6 nA.
    04:50 - Polarization drops to 43%.
    05:41 - Due to continues decrease of target polarization (below 40%), we
    stopped the beam to improve.
    06:03 - Beam back on. Polarization -53.7%. 
    06:08 - Polarization drop to -43.7% in 5 min. 
    06:08 - Decreasing beam current to improve polarization.
    06:24 - Spot++ looks fine.
    06:48 - IHWP now OUT.
    06:58 - Beam position epics variables on strip chart are reading 0.
    spot++ looks fine.Maybe low current is responsible?
    07:21 - Spot++ looks fine.
    07:43 - Polarization now is -52.3%.